Banyan Co.’s common stock currently sells for $41.00 per share. The growth rate is a constant 13.5%, and the company has an expected dividend yield of 4%. The expected long-run dividend payout ratio is 25%, and the expected return on equity (ROE) is 18%. New stock can be sold to the public at the current price, but a flotation cost of 10% would be incurred. What would be the cost of new equity? Round your answer to two decimal places. Do not round your intermediate calculations.

%

please Help

Thank you!
Answer #1

Current Stock price = $41.00
Constant growth rate= 13.5%
Expected dividend yield = 4%
Expected long-run dividend payout ratio = 25%
Expected return on equity= 18%
Flotation cost= 10%
Cost of New Equity =?

Growth rate = (1-Dividend payout ratio)*(ROE)
= (1-25%)*(18%)=0.135

Cost of new equity= [(Expected Dividend)/(Current Stock price*(1-Flotation cost))] + Constant growth rate

Dividend yield = (Expected Dividend)/Current Stock price
4%= (Expected Dividend)/41
=>Expected Dividend=4%*41=1.64

Cost of new equity = [(Expected Dividend)/(Current Stock price*(1-Flotation cost))] + Constant growth rate
Cost of new equity = [(1.64)/(41*(1-10%))] + 13.5%
=(1.64/36.9) + 13.5%
=0.179444444 or 17.94% (Rounded to two decimal places)

So, the cost of new equity is 17.94%